Bank robbers brandish firearms to threaten their victims in 80% of the incidents. An announcement that six bank robbers are taking place is being broadcast. What is the probability that a firearm is being used in atleast one of the robberies?
Answer by stanbon(75887)

Bank robbers brandish firearms to threaten their victims in 80% of the incidents. An announcement that six bank robbers are taking place is being broadcast. What is the probability that a firearm is being used in atleast one of the robberies?
---------
Binomial Problem with n = 6 and p(gun) = 0.8
----
P(1<= x <=6) = 1 - P(x = 0)
---
= 1 - 0.2^6
---
= 1 - 0.000064
================
= 0.9999
===========================
Cheers,
Stan H.
===============